Reid has “n” apples. When he is given 30 more he has four times as many as he started with. Find the value of “n”
Answer by ikleyn(52778)

.

    n + 30 = 4n

    30     = 4n - n

    30     = 3n

     n     = 30/3 = 10.    ANSWER

Solved.